OpenStudy (anonymous):

Which method would be easiest way to solve this equation: x+2y=-5 -x+y=-4 *and can you answer it, thanks!

OpenStudy (anonymous):

simply add them together

OpenStudy (anonymous):

once you add them you get: 3y=-9, now to solve for y divide both sides by 3 and you get y=-3

OpenStudy (anonymous):

and that method is called elimination method

OpenStudy (anonymous):

now substitute -3 into the first equation x+2y=-5, like so x+2(-3)=-5

OpenStudy (anonymous):

now solve x+2(-3)=-5 for x

OpenStudy (anonymous):

x = 1

OpenStudy (anonymous):

your correct

OpenStudy (anonymous):

thanks for your help!